### v2

@@ 4 Conclusions
For explicit models, the rules governing the spatial movement of viruses appear to affect the infection more significantly than the rules governing tumour cell movement. Nonetheless, cell movement does still retain its influence on tumour invasion, especially when movement is constrained: only pressure-driven equations describe a slower invasion for low tumour cell numbers. This is a relevant aspect in the overall success or failure of the therapy and bears clinical consequences that are important when analysing patient’s results. Overall, undirected movement (with or without viral diffusion) and pressure-driven movement with viral diffusion exhibit similar behaviour in regards to the spread of the infection throughout a tumour. The main difference among all models is the speed of the infections and, when this is crucial for applications, our result suggest that the viral population should be openly included in the formulation. An additional difference that has therapeutic value is the appearance of oscillations between viral and tumour populations, which is not a feature shared by all models. For instance, a pressure-driven tumour not explicitly infected by a viral population (or with a low viral diffusion coefficient) exhibits a very different behaviour than a model with unhindered tumour diffusion. Notably, the presence of a viral population can also lead to persistent oscillations, which, as we have seen, can significantly affect the overall outcome in some cases. This could be very valuable when considering, for example, the effect of viral vectors on outcomes for some chronic blood tumours, which have been known for many years to exhibit semi-regular oscillatory regimes [CHIKKAPPA1980Blood, RODRIGUEZ1976AMJ, Vodopick1972NEJM].
@@
In conclusion, it is important to remark that, for pressure-driven models, cells’ displacement alone cannot spread viral infection and central infections remain localised. As previously observed, even wide infections may be stopped by stochastic events [morselli23], but the situation significantly changes when viral particles are allowed to freely move around the tumour and cell-to-cell contact becomes less relevant for the infection. In that case, our theoretical and computational results have explained that this infection is qualitatively similar to the case of undirected cell movement. We can conclude that pressure-driven movement allows us to describe both the situations of constrained and unconstrained cellular and viral movements, while such constraints cannot be modelled with undirected movement. The increased realism of pressure-driven movement, though, comes at the cost of a lower degree of theoretical understanding of the models.
Finally, a couple of possible extensions of this work are worth being mentioned. Firstly, it would be interesting to include actual immunotherapies and other contemporary clinical therapies in the explicit models presented here. This could allow to extend the theoretical and computational predictions we have discussed and compare them with more realistic scenarios. The evaluation of the role of spatial dependency for viral infection could also particularly interesting. For example, looking at how tumour shape influences the outcome of virotherapy or where and how the way viral load is injected can give rise to different infection speeds are both valuable ideas for future works.
@@ A.1 Uninfected cells
$\partial_{t}u=D\partial_{xx}^{2}u+p\Bigl(1-\frac{u}{K}\Bigr)u$
@@
$C=\frac{u}{K},\quad\hat{t}=pt,\quad\hat{x}=x\sqrt{\frac{p}{D}}$
@@
$\partial_{t}u=\frac{D}{K}\partial_{x}(u\partial_{x}u)+p\Bigl(1-\frac{u}{K}\Bigr)u$
@@
Results analogous to the previous case still hold, but this time the minimum speed is $c^{*}\coloneqq 1/\sqrt{2}$ [aronson80, newman80]. Its dimensional value is $\sqrt{Dp/2}$. It is straightforward to check that we recover the same speed value as in Eq. (A.1) by multiplying the diffusion coefficient by a factor of $8$: this is indeed the procedure that we follow whenever we compare models with different rules of movement, so that we avoid any bias due to the different propagation speed.
@@
$C=\frac{u}{\bar{u}},\quad\hat{t}=pt,\quad\hat{x}=x\sqrt{\frac{pK}{D\bar{u}}}$
@@
$\sqrt{\frac{Dp}{2}\ \frac{\bar{u}}{K}}$ (A.3)
@@ A.2 Infected cells and virus
Invasion into an uninfected region at carrying capacity (following Ref. [baabdulla23])
@@
$D_{c}=\frac{D}{D_{v}},\quad a=\frac{q}{p},\quad\theta=\frac{\alpha\beta q}{p^{2}},\quad\gamma=\frac{q_{v}}{p}$
@@
$\Phi(\lambda,c)\coloneqq D_{c}\lambda^{4}-c(D_{i}+1)\lambda^{3}+(c^{2}-D_{c}\gamma-a)\lambda^{2}+c(\gamma+a)\lambda+a\gamma-\theta=0$
@@
$c^{\prime}(\lambda)=\frac{\Psi(\lambda;c)}{2c\lambda^{2}-(D_{i}+1)\lambda^{3}+(\gamma+a)\lambda}$
@@
$\Psi(\lambda,c)\coloneqq 2\lambda c^{2}+(\gamma+a-3(D_{c}+1)\lambda^{2})c+4D_{i}\lambda^{3}-2(D_{c}\gamma+a)\lambda$
Hence the minimum wave speed $c^{*}$ arises at the intersection of the two manifolds $\{\Phi(\lambda,c)=0\}\cap\{\Psi(\lambda,c)=0\}$. The value of $c^{*}$ can be found numerically, using the built-in Matlab function fsolve. The dimensional value of the speed is then $c^{*}\sqrt{D_{v}p}$.
@@
Figure 11: Dimensional minimum speed for different values of $\bar{u}$ compared to the speed of the uninfected travelling wave $2\sqrt{Dp}$. The parameters employed are the ones given in Table 1, with both values of $\alpha$ and $q_{v}$. The higher value of $D_{v}$ is the reference one, while the lower value is $D_{v}=10^{-5}\;$mm2/h.
@@
$\tilde{\Phi}(\lambda,c)\coloneqq D_{c}\lambda^{4}-c(D_{i}+1)\lambda^{3}+(c^{2}-D_{c}\gamma-a)\lambda^{2}+c(\gamma+a)\lambda+a\gamma-\theta\bar{u}=0$

### v3

@@ 4 Conclusions
For explicit models, the rules governing the spatial movement of viruses appear to affect the infection more significantly than the rules governing tumour cell movement. Nonetheless, cell movement does still retain its influence on tumour invasion, especially when movement is constrained: only pressure-driven equations describe a slower invasion for low tumour cell numbers. This is a relevant aspect in the overall success or failure of the therapy and bears clinical consequences that are important when analysing patient’s results. Overall, undirected movement (with or without viral diffusion) and pressure-driven movement with viral diffusion exhibit similar behaviour in regards to the spread of the infection throughout a tumour. The main difference among all models is the speed of the infections and, when this is crucial for applications, our result suggest that the viral population should be openly included in the formulation. An additional difference that has therapeutic value is the appearance of oscillations between viral and tumour populations, which is not a feature shared by all models. For instance, a pressure-driven tumour not explicitly infected by a viral population (or with a low viral diffusion coefficient) exhibits a very different behaviour than a model with unhindered tumour diffusion. Notably, the presence of a viral population can also lead to persistent oscillations, which, as we have seen, can significantly affect the overall outcome in some cases. This could be very valuable when considering, for example, the effect of viral vectors on outcomes for some chronic blood tumours, which have been known for many years to exhibit semi-regular oscillatory regimes [16, 64, 73].
@@
In conclusion, it is important to remark that, for pressure-driven models, cells’ displacement alone cannot spread viral infection and central infections remain localised. As previously observed, even wide infections may be stopped by stochastic events [53], but the situation significantly changes when viral particles are allowed to freely move around the tumour and cell-to-cell contact becomes less relevant for the infection. In that case, our theoretical and computational results have explained that this infection is qualitatively similar to the case of undirected cell movement. We can conclude that pressure-driven movement allows us to describe both the situations of constrained and unconstrained cellular and viral movements, while such constraints cannot be modelled with undirected movement. The increased realism of pressure-driven movement, though, comes at the cost of a lower degree of theoretical understanding of the models.
Finally, a couple of possible extensions of this work are worth being mentioned. Firstly, it would be interesting to include actual immunotherapies and other contemporary clinical therapies in the explicit models presented here. This could allow to extend the theoretical and computational predictions we have discussed and compare them with more realistic scenarios. The evaluation of the role of spatial dependency for viral infection could also be particularly interesting. For example, looking at how tumour shape influences the outcome of virotherapy or where and how the way viral load is injected can give rise to different infection speeds are both valuable ideas for future works.
@@ A.1 Uninfected cells
$\partial_{t}u=D_{U}\partial_{xx}^{2}u+p\Bigl(1-\frac{u}{K}\Bigr)u$
@@
$C=\frac{u}{K},\quad\hat{t}=pt,\quad\hat{x}=x\sqrt{\frac{p}{D_{U}}}$
@@
$\partial_{t}u=\frac{D_{P}}{K}\partial_{x}(u\partial_{x}u)+p\Bigl(1-\frac{u}{K}\Bigr)u$
@@
Results analogous to the previous case still hold, but this time the minimum speed is $c^{*}\coloneqq 1/\sqrt{2}$ [3, 55]. Its dimensional value is $\sqrt{D_{P}p/2}$. It is straightforward to check that we recover the same speed value as in Eq. (A.1) by multiplying the diffusion coefficient by a factor of $8$: this is indeed the procedure that we follow whenever we compare models with different rules of movement, so that we avoid any bias due to the different propagation speed.
@@
$C=\frac{u}{\bar{u}},\quad\hat{t}=pt,\quad\hat{x}=x\sqrt{\frac{pK}{D_{P}\bar{u}}}$
@@
$\sqrt{\frac{D_{P}p}{2}\ \frac{\bar{u}}{K}}$ (A.3)
@@ A.2 Infected cells and virus
Invasion into an uninfected region at carrying capacity (following Ref. [4])
Figure 12: Theoretical speed of the travelling wave of viral infection invading an uninfected tumour at carrying capacity for different parameter values compared to the speed of uninfected front (black dotted line) and the speed of infection for Eq. (U2) (cyan dashed line). In panel (a), the value of $\tilde{\beta}$ is not constant, so the comparison with Eq. (U2) is not particularly meaningful. Observe that for low values of $\alpha$ there is no travelling wave, as $i^{*}$ is negative and the equilibrium $(u^{*},i^{*})$ is unstable (see also Fig. 6).
@@
$D_{c}=\frac{D_{U}}{D_{v}},\quad a=\frac{q}{p},\quad\theta=\frac{\alpha\beta q}{p^{2}},\quad\gamma=\frac{q_{v}}{p}$
@@
$\Phi(\lambda,c)\coloneqq D_{c}\lambda^{4}-c(D_{c}+1)\lambda^{3}+(c^{2}-D_{c}\gamma-a)\lambda^{2}+c(\gamma+a)\lambda+a\gamma-\theta=0$
@@
$c^{\prime}(\lambda)=\frac{\Psi(\lambda;c)}{2c\lambda^{2}-(D_{c}+1)\lambda^{3}+(\gamma+a)\lambda}$
@@
$\Psi(\lambda,c)\coloneqq 2\lambda c^{2}+(\gamma+a-3(D_{c}+1)\lambda^{2})c+4D_{c}\lambda^{3}-2(D_{c}\gamma+a)\lambda$
Hence the minimum wave speed $c^{*}$ arises at the intersection of the two manifolds $\{\Phi(\lambda,c)=0\}\cap\{\Psi(\lambda,c)=0\}$. The value of $c^{*}$ can be found numerically, using the built-in Matlab function fsolve. The dimensional value of the speed is then $c^{*}\sqrt{D_{v}p}$, which is shown in Fig. 12 for different parameter values. Observe that higher values of $\alpha$ and lower values of $q_{v}$ result in a higher value of $\tilde{\beta}$, corresponding to a more effective infection (in line with the results of Section 3.2). Similarly, the infection propagates faster for high values of $D_{v}$, although equilibrium values are not affected.
Figure 13: Dimensional minimum speed for different values of $\bar{u}$ compared to the speed of the uninfected travelling wave $2\sqrt{D_{U}p}$. The parameters employed are the ones given in Table 1, with both values of $\alpha$ and $q_{v}$. The higher value of $D_{v}$ is the reference one, while the lower value is $D_{v}=10^{-5}\;$mm2/h.
@@
$\tilde{\Phi}(\lambda,c)\coloneqq D_{c}\lambda^{4}-c(D_{c}+1)\lambda^{3}+(c^{2}-D_{c}\gamma-a)\lambda^{2}+c(\gamma+a)\lambda+a\gamma-\theta\bar{u}=0$
@@
Supplementary information
The supplementary material is available in the online version of the published paper, as well as at the link https://doi.org/10.5281/zenodo.18877438.